OpenStudy (anonymous):

What is the solution to this problem...Pipe A can fill in 20 Minutes, Pipe B in 30 Minutes; pipe C can empty the same in 40 Minutes. If all of them work together, find the time taken to fill the tank.

OpenStudy (anonymous):

First of all, think for just 1 minute: \[\frac{ 1 }{ 20 }+\frac{ 1 }{ 30 }-\frac{ 1 }{ 40 }=\frac{ 7 }{ 120 }\] So all pipes together fill 7/120 of the tank in 1 minute. In order to find the time taken to fill the tank: \[\frac{ 120 }{ 7 } \approx 17.14\] in terms of minutes.